OpenStudy (jordanloveangel):

WHich is the graph of y= 1/2x?

ganeshie8 (ganeshie8):

\[\large \rm y = \dfrac{1}{2}x\]

OpenStudy (some.random.cool.kid):

probably graph A bye

OpenStudy (anonymous):

a

OpenStudy (jordanloveangel):

yeah i was thinking A 2

ganeshie8 (ganeshie8):

thats a very good wrong guess

OpenStudy (anonymous):

its c

OpenStudy (jordanloveangel):

realy?

ganeshie8 (ganeshie8):

slope = 1/2 rise = 1 run = 2

OpenStudy (anonymous):

Its C

ganeshie8 (ganeshie8):

that means the graphs goes UP 1 unit for every 2 units RIGHT yes ?

OpenStudy (jordanloveangel):

y = (1/2)^x is an exponential function where x is the exponent 1° case when x approaches negative infinity Note that y = (1/2)^x = 2^(-x) so when x -> (-inf.), function approaches y = 2^[-(-inf)] = 2^(+inf) and lim(y) {x -> -inf} = +inf When x decreases to -inf., y grows to +inf 2° case when x approaches positive infinity We can write y = (1/2)^x = f(x)/g(x) where f(x)=1 and g(x)=2^x and then use following rule lim[f(x)/g(x)] {x -> inf} = lim f(x) {x -> inf} / lim g(x) {x -> inf} therefore lim[(1/2)^x] {x -> inf} = lim[1/(2^x)] {x -> inf} = lim (1){x -> inf} / lim (2^x){x -> inf} lim (1){x -> inf} = 1 lim (2^x){x -> inf} = inf lim[(1/2)^x] {x -> inf} = 1 / inf. = 0 The limit of y as x approaches +infinity is 0. That means horizontal line y=0 is horizontal asymptote of function y = (1/2)^x. 3° intercept with y-axis For x=0, y = (1/2)^0 = 1 Graph of this function passes through (0,1).

ganeshie8 (ganeshie8):

thats a different question y = (1/2)x is not same as y = (1/2)^x

ganeshie8 (ganeshie8):

btw C is the right graph

OpenStudy (some.random.cool.kid):

im not sure its C

OpenStudy (some.random.cool.kid):

i would have to go with a

OpenStudy (some.random.cool.kid):

because theres a direct split in graph A

ganeshie8 (ganeshie8):

A is y = 2x @some.random.cool.kid

OpenStudy (jordanloveangel):

im confused

OpenStudy (jordanloveangel):

@ganeshie8 the right answer is C

ganeshie8 (ganeshie8):

yes

OpenStudy (jordanloveangel):

thx